# Prompt-Counterfactual Explanations for Generative AI System Behavior

### v1

@@ Abstract
As generative AI systems become integrated into real-world applications, organizations increasingly need to be able to understand and interpret their behavior. In particular, decision-makers need to understand what causes generative AI systems to exhibit specific output characteristics. Within this general topic, this paper examines a key question: what is it about the input—the prompt—that causes an LLM-based generative AI system to produce output that exhibits specific characteristics, such as toxicity, negative sentiment, or political bias. To examine this question, we adapt a common technique from the Explainable AI literature: counterfactual explanations. We explain why traditional counterfactual explanations cannot be applied directly to generative AI systems, due to several differences in how generative AI systems function. We then propose a flexible framework that adapts counterfactual explanations to non-deterministic, generative AI systems in scenarios where downstream classifiers can reveal key characteristics of their outputs. Based on this framework, we introduce an algorithm for generating prompt-counterfactual explanations (PCEs). Finally, we demonstrate the production of counterfactual explanations for generative AI systems with three case studies, examining different output characteristics (viz., political leaning, toxicity, and sentiment). The case studies further show that PCEs can streamline prompt engineering to suppress undesirable output characteristics and can enhance red-teaming efforts to uncover additional prompts that elicit undesirable outputs. Ultimately, this work lays a foundation for prompt-focused interpretability in generative AI: a capability that will become indispensable as these models are entrusted with higher-stakes tasks and subject to emerging regulatory requirements for transparency and accountability.
Keywords: Explainable AI, Generative AI, LLMs, Counterfactual Explanations
@@ 1 Introduction
As firms work to integrate large-scale generative AI systems across functions, they increasingly need to be able to understand not just what the systems generate, but also why they generate it. Generative AI systems are now being used in sales, marketing, operations, research, recommender systems, healthcare, education, and many more functions (raj2023analyzing; kasneci2023chatgpt; singhal2023large; gomez2023confederacy; friedman2023leveraging). However, generative AI system outputs are complex, typically unstructured, and often non-deterministic, which makes them quite challenging to monitor, evaluate, analyze, and explain.
Generative AI refers to a broad class of systems that produce text, images, audio, or other content based on learned distributions (feuerriegel2024generative). In this paper, we will focus on AI systems based on large language models (LLMs) such as GPT-4 (achiam2023gpt), Claude 3.7 Sonnet (anthropic2024claude), or Llama 3.3 (grattafiori2024llama), which currently represent the most widely applied subclass of generative AI systems. [^1] We use publicly available models on HuggingFace model library https://huggingface.co/models as a proxy. As of Dec 7th, 2025, there are 300,930 out of 2,263,421 models that are text generation models. Despite our specific focus on LLMs, this methodology can be applied to other generative modalities.
@@
In this paper, we propose a framework for adapting and applying counterfactual (CF) explanations (martens2014explaining; wachter2017counterfactual; fernandez_explaining) to generative AI systems. Counterfactual explanations, a popular method for explaining the behavior of traditional predictive systems, identify (minimal) changes to the input that would alter the AI system’s prediction, decision, or action. However, traditional CF explanations cannot be applied directly to generative AI systems for several reasons, requiring rethinking and extending the notion of CF explanations. Except in very restrictive settings, generative AI systems do not produce discrete predictions, decisions, or actions (hereafter, decisions). Instead, they generate complex outputs; for example, LLMs produce open-ended text. In addition, the outputs of the generative AI system are very often stochastic. Even with an identical input prompt, the outputs can vary from run to run.
For this paper, we focus on explaining a specific class of AI system behaviors: whether the output exhibits a certain characteristic (often undesirable). Behaviors in this class include things like: did the system produce hate speech? Is something in the system’s output fabricated? Does the system’s output violate company policy or otherwise require content moderation action? Is the system’s output biased in a certain way? Does the system’s output exhibit very negative sentiment (hartvigsen2022toxigen; abid2021persistent; perez2022red; huang2025survey)? Other applications may require the identification of task-specific characteristics. For example, in personalized recommender systems, a firm may want to evaluate whether the generated suggestions align with a user’s preferences and therefore measure elements such as diversity, novelty and bias in the output (chen2024large). For a deployed chat system, we may want to assess whether it exhibits traits analogous to empathy, or analyze its generated conversation with respect to the Big Five personality traits (concannon2024measuring; liu2025illusion; salecha2024large). We will refer to this sort of behavior characteristic as downstream classification; ideally such characteristics can be estimated well via classifiers downstream from the generated output. [^2] Here, we assume that such a sufficient classifier exists; whether a particular generated output characteristic truly can be estimated well will depend on the characteristic. This paper focuses on explaining what about the input leads generative AI outputs to exhibit a particular behavior as revealed by downstream classification.
@@
The rest of the paper is structured as follows. In Section 2, we contextualize our contribution with prior work on Explainable AI in general and specifically for generative AI systems. We present the problem statement and the setup for generative AI PCEs in Section 3. Here, we discuss how to adapt counterfactual explanations to deal with generative AI’s non-determinism, by focusing on aggregate properties of the model outputs. In Section 4, we present a straightforward, first PCE algorithm for the input/output behavior of the generative AI system. Section 5 presents results across three illustrative case studies, highlighting scenarios where PCEs could be useful. Specifically, the case studies focus on three different downstream classifications: political leaning detection, toxicity prediction, and sentiment classification. We end the paper in Section 6 with a discussion and report limitations and avenues for future research.
@@ 2.1 Explainable AI and Counterfactual Explanations
As AI systems become increasingly integrated into business processes, the demand for transparency in these systems grows. In addition to the AI system being effective, it is often important that developers, users, and other stakeholders receive insight into the why and how behind the behavior of the system. These needs for transparency are described in detail by martens2014explaining.
The research field addressing techniques to provide such transparency has become known as Explainable AI (XAI). XAI techniques aim to make the internal mechanisms or the reasons for the input/output behavior of AI systems more comprehensible to humans (gunning2019xai). More specifically, XAI methods examine a variety of different aspects of AI systems, and researchers and practitioners need to be precise about what sort of AI explanation they are talking about (martens2025beware). The vast majority of XAI techniques focus on one of the following: building interpretable models, understanding the inner workings of complex models, identifying features that affect model scoring, or identifying features or parts of inputs that lead to a certain system behavior. This paper focuses on the last of those.
AI explanations serve a variety of purposes. They can help build user or management trust, assist model developers in debugging or model improvement, uncover hidden biases or fairness concerns, satisfy regulatory or legal requirements for transparency, improve relations with customers affected by AI systems actions, etc. (ferrario2022explainability; goethals2024precof; martens2014explaining; vermeire2022explainable; wachter2017counterfactual). XAI methods have been developed to explain both global behavior (how a model behaves across the input space), and local behavior (why a model made a specific prediction or decision for a given instance) (martens2014explaining; guidotti2018survey; molnar2020interpretable).
This paper focuses on explanations for local (instance-specific) input/output behavior of generative AI systems. We adapt a well-known instance-level explanation technique, namely counterfactual explanations (fernandez_explaining; martens2014explaining; wachter2017counterfactual). Counterfactual explanations aim to identify (minimal) changes to an input instance that would result in a different decision outcome. For example, in the context of credit scoring, a possible counterfactual explanation could be ‘If the loan amount had been \$16,000 lower, you would have received the loan’ (fernandez_explaining).
@@ 2.2 Explanations for Generative AI systems
LLMs like ChatGPT have been criticized for being opaque, black boxes (das2025security; zhao2024explainability). The internal workings of LLMs remain largely inscrutable, even to their developers, and foundational model providers rarely provide explanations for their systems’ behavior. This is particularly problematic given the systems’ remarkable performance and attraction to client organizations, as their opaqueness poses challenges for safety, accountability, human trust, and even debugging functional systems built based on foundational models. Explainable AI techniques are essential tools to foster a deeper understanding of LLM behavior and to mitigate the risks associated with their deployment (das2025security; ferdaus2024towards).
@@
A significant portion of the literature on LLM interpretability focuses on model-centric approaches, specifically, approaches that reference the internals of the model. For example, representation surgery involves directly intervening in a model’s latent space to alter internal representations or isolate specific features (ravfogelgumbel). More specifically, ravfogelgumbel study the idea of counterfactuals in the representation space, raising the question: What would the model have generated if its internal representation had undergone a specific intervention? [^4] As noted above, the term “counterfactual” is overloaded even in the XAI space; ravfogelgumbel employ a different notion of a counterfactual than the notion we adapt from the XAI literature. (To our knowledge the latter usage was the original use of the term for XAI.) Other methods, such as circuit analysis, attempt to decompose computations into interpretable submodules or “circuits” to trace how information flows through the network (tang2023large). Attention-based explanations focus on analyzing the knowledge encoded in attention weights (zhao2024explainability). These internalist approaches require deep technical expertise, are difficult to transfer across different model architectures, and may or may not produce an “explanation” for the AI system’s behavior that is actually interpretable to stakeholders.
@@
Our work fits into the body of research that adopts a prompt-centric perspective to analyzing LLM behavior, examining how variation in inputs, rather than model internals or training data, shape the behavior of LLMs. Numerous studies show that LLMs are highly prompt-sensitive, with even minor changes significantly affecting outputs (anagnostidis2024susceptible; elazar2021measuring; rauba2024quantifying). rauba2024quantifying highlight the obstacle of disentangling meaningful changes in the output from the inherent stochasticity of LLM outputs, and propose a statistical framework to reformulate LLM perturbation analysis as a frequentist hypothesis problem. mohammadi2024explaining leverage Shapley values to expose the “token noise” effect, where seemingly unimportant tokens exert a disproportionate influence on model output. Tools like Polyjuice generate diverse, plausible counterfactual prompts to support this line of inquiry (wu2021polyjuice). Other studies investigate the performance of using LLMs to generate counterfactual explanations themselves (li2024prompting; youssef2024llms; mayne2025llms).
@@
Our work examines how counterfactual explanations can be meaningfully applied to non-deterministic generative AI systems, when the focal characteristics of generative outputs can be revealed using downstream classification. To the best of our knowledge, we are the first to apply counterfactual explanations to full text continuations, whereas earlier work primarily focused on providing counterfactual explanations in contexts where generative AI systems were used to make explicit decisions, for example labeling (chittimalla2025explainable; mayne2025llms; sarkar2024explaining).
@@ 3 Problem formulation
Explainable AI can be confusing because there are so many different ways explanations might be defined. Here we follow and adapt previous work on counterfactual explanations for AI decisions (fernandez_explaining). However, there are several non-trivial challenges to doing so.
@@ 3.1 Four key challenges
    The majority of existing counterfactual methods assume feature-set input, and the methods operate on these sets of features; in particular, the ordering of the input is ignored. Although the original counterfactual explanation method operated on text as input (martens2014explaining), it essentially assumed a token-set representation (bag-of-words, tf-idf, etc.), without considering the sequential and linguistic structure of the text. While a token-set explanation method seems like a great place to start, we would not want to use it to define explanations for LLM-based systems, as the linguistic structure of the text may be critical. An exception to simple feature-set input in prior work is CFEs for image classification, where “pieces” (e.g., regions) of the image can be defined in different ways and then operated on to produce counterfactuals (vermeire2022explainable); we use this idea as inspiration for our adaptation to LLMs.
@@
    The set of counterfactual explanations that will be produced depends on how one searches through the space of collections of “pieces” of the input to remove (see above), and with what (if anything) one replaces them. How this replacement works is a crucial aspect of formulating explanations that will be effective for a particular problem (fernandez_explaining), and as Fernández et al. explained, the choice of replacement should depend on the specifics of the particular explanation context. Explanation methods should be flexible enough to allow for different choices. For LLM-based generative systems, we can envision many different replacement options, including dropping pieces of the prompt, masking them, replacing them with synonymous phrases, and so on. In addition, the structure of the text may lead to more intelligent search strategies for choosing combinations of “pieces” of the input to replace, which could also lead to significant speed-ups for the algorithm. A straightforward example would be to prefer combining adjacent input pieces (sequential tokens, phrases, sentences). One can envision more sophisticated techniques taking into account the vast amount of syntactic and semantic analysis provided by (computational) linguistics.
@@
    Finally, unlike prior applications of CF explanations, generative AI systems do not produce a fixed, deterministic output from a given input (atil2024non). This substantially complicates matters. CFEs ask some variant of “what can I change in the input to get a focal change in the output?” But just running the system again with the same input might produce the focal change in the output! Therefore, the question needs to be generalized to be probabilistic or otherwise distributional. For example, a PCE algorithm could ask about (changes to) the probability that a particular input will produce an output with/without the focal characteristic.
@@
In order to change the input, we mainly experiment with a straightforward approach that involves replacing the chosen segments of the prompt with a textual “masking” token (underscores). [^6] Using explicit token masking produced comparable results. There is a difference between the “replacement” strategy used to generate explanations, and a replacement strategy used to produce suitable output based on what the explanations revealed. The PCEs are an analysis tool. There always will still be the question: what should we do based on the results? [^7] The counterfactual explanations literature sometimes confounds the questions of what in its input caused the system to give a particular output—the primary CFE question—and what can be done in the real world based on the CFE answer. In the case studies, when looking into what we might do based on a PCE, we demonstrate alternatives to simply removing segments, such as replacing them with paraphrases when that makes sense for the application.
@@
To illustrate, let us say that the characteristic we are interested in is the sentiment of our LLM’s generated output. Imagine that the generative system receives the prompt “Create a horrible short story about a dog” and our characteristic classifier is a sentiment classifier that classifies the polarity and magnitude of the sentiment of the generated output. Due to the word “horrible,” it would be natural to expect that the output for this prompt would have negative sentiment polarity. Figure 2(a) shows the distribution over the sentiment of the outputs over 100 different generations based on this one prompt. The outputs are often, but not always, classified as having negative sentiment.
(a) Polarity distribution of the output of the generative AI system after the prompt “Create a horrible short story about a dog” was run 100 times. We see that the sentiment of the generated output varies widely, although it very often is very negative. The vertical dashed line shows the average of the sentiment scores across the 100 runs.
(b) Polarity distribution of the output of the generative AI system after the prompt “Create a short story about a dog” was run 100 times. We see that removing the word “horrible,” identified by the PCE, shifts the sentiment to be generally much more positive (the dashed line shows the average sentiment score). Notably, now there are very few cases where the generated outputs have very negative sentiment.
@@
Of course there is not a single right answer to either of these questions; we need to define what we care about for the particular task at hand. [^9] This is not a new observation; it is consistent with prior literature on counterfactual explanations for AI systems fernandez_explaining. Thus, a framework for creating PCEs for generative AI systems needs to be flexible. Do we care about the mean prediction probability of negative sentiment over all the samples? The median? Where do we want to set the threshold in any of these cases in order to say that the behavior has changed? How many runs should we look at? The definition of a PCE will depend on the user’s needs. For example, in one setting, we might care that the sentiment the system produces is generally positive. In another setting, we might be extremely averse to negative sentiment, and therefore specify that the likelihood of negative sentiment must be very small.
To illustrate, let’s say we decide to use the average (mean) of the characteristic scores (from the downstream classifier) across the empirical distribution, and we want this average to be positive. Based on this choice, we can define a PCE as a minimal set of “parts” of the input prompt that, if they were not present, would move the average of the prediction distribution from below to above the threshold (in this case 0). Our explanation algorithm (explained in detail in Section 4) for the dog story example leads to the not-surprising explanation: “If the word horrible were not part of the input prompt, the average sentiment of the generated output would be positive.” The resultant prediction distribution, with “horrible” removed, is also depicted in Figure 2(b).
Importantly, as mentioned above, this tells us what is it about the prompt that resulted in the sentiment of the output to be negative. It does not (necessarily) tell us what to do about it. That is a different problem from creating the PCE. It is unclear for this example whether the system could create a horrible story about a dog that did not have negative sentiment. In our case studies, we illustrate several possible ways to act on this information, such as by paraphrasing the “offending” parts of the input identified by the PCE.

### v2

@@ Abstract
As generative AI systems become integrated into real-world applications, organizations increasingly need to be able to understand and interpret their behavior. This paper examines a key question: what is it about the input—the prompt—that causes an LLM-based generative AI system to produce output that exhibits specific characteristics. We adapt a common technique from the Explainable AI literature: counterfactual explanations. We explain why traditional counterfactual explanations cannot be applied directly to generative AI systems, due to four challenges, and then propose a flexible framework that adapts counterfactual explanations to non-deterministic, generative AI systems in scenarios where downstream classifiers can reveal key characteristics of their outputs. Based on this framework, we introduce an algorithm for generating prompt-counterfactual explanations (PCEs). We demonstrate the production of counterfactual explanations with three case studies, exhibiting different output characteristics (viz., political leaning, toxicity, and sentiment) and different characteristic distributions. The case studies demonstrate the production of PCEs, addressing the four challenges, and suggest that PCEs could streamline prompt engineering to suppress undesirable output characteristics and enhance red-teaming efforts. Ultimately, this work lays a foundation for prompt-focused interpretability in generative AI, a necessary capability as these models are applied to higher-stakes tasks with requirements for transparency and accountability.
Keywords: Explainable AI, Generative AI, LLMs, Counterfactual Explanations, Transparency
@@ 1 Introduction
As firms work to integrate large-scale generative AI systems across functions, they increasingly need to be able to understand not just what the systems generate, but also why they generate it. Generative AI systems are now being used in sales, marketing, operations, research, recommender systems, healthcare, education, and many more functions (Raj et al., 2023; Kasneci et al., 2023; Singhal et al., 2023; Gómez-Rodríguez et al., 2023; Friedman et al., 2023). However, generative AI system outputs are complex, typically unstructured, and often non-deterministic, which makes them quite challenging to monitor, evaluate, analyze, and explain.
Generative AI refers to a broad class of systems that produce text, images, audio, or other content based on learned distributions (Feuerriegel et al., 2024). In this paper, we will focus on AI systems based on large language models (LLMs), which currently represent the most widely applied subclass of generative AI systems. [^1] We use publicly available models from the HuggingFace model library https://huggingface.co/models for our experiments. As of Dec 7th, 2025, 300,930 out of 2,263,421 HuggingFace models that text generation models. Despite our specific focus on LLMs, this methodology can be applied to other generative modalities.
@@
In this paper, we propose a framework for adapting and applying counterfactual (CF) explanations (Martens and Provost, 2014; Wachter et al., 2017; Fernández-Loría et al., 2022) to generative AI systems. Counterfactual explanations, a popular method for explaining the behavior of traditional predictive systems, identify (minimal) changes to the input that would alter the AI system’s prediction, decision, or action. However, traditional CF explanations cannot be applied directly to generative AI systems for several reasons, requiring rethinking and extending the notion of CF explanations. Except in very restrictive settings, generative AI systems do not produce discrete predictions, decisions, or actions (hereafter, decisions). Instead, they generate complex outputs; for example, LLMs produce open-ended text. In addition, the outputs of the generative AI system are very often stochastic. Even with an identical input prompt, the outputs can vary from run to run.
For this paper, we focus on explaining a specific class of AI system behaviors: whether the output exhibits a certain characteristic (often undesirable). Behaviors in this class include things like: did the system produce hate speech? Is something in the system’s output fabricated? Does the system’s output violate company policy or otherwise require content moderation action? Is the system’s output biased in a certain way? Does the system’s output exhibit very negative sentiment (Hartvigsen et al., 2022; Abid et al., 2021; Perez et al., 2022; Huang et al., 2025)? Other applications may require the identification of task-specific characteristics. For example, in personalized recommender systems, a firm may want to evaluate whether the generated suggestions align with a user’s preferences and therefore measure elements such as diversity, novelty and bias in the output (Chen et al., 2024). For a deployed chat system, we may want to assess whether it exhibits traits analogous to empathy, or analyze its generated conversation with respect to the Big Five personality traits (Concannon and Tomalin, 2024; Liu et al., 2025; Salecha et al., 2024). We will refer to this sort of behavior characteristic as downstream classification; ideally such characteristics can be estimated well via classifiers downstream from the generated output. [^2] For this paper, we assume that such a sufficient classifier exists; whether a particular generated output characteristic truly can be estimated well will depend on the characteristic. This paper focuses on explaining what about the input leads generative AI outputs to exhibit a particular behavior as revealed by downstream classification.
@@
The rest of the paper is structured as follows. In Section 2, we contextualize our contribution with prior work on Explainable AI in general and specifically for generative AI systems. We present the problem statement and the setup for generative AI PCEs in Section 3. Here, we discuss how to adapt counterfactual explanations to deal with generative AI’s non-determinism, by focusing on aggregate properties of the model outputs. In Section 4, we present a straightforward, first PCE algorithm for the input/output behavior of the generative AI system. Section 5 presents results across three illustrative case studies, highlighting scenarios where PCEs could be useful. Specifically, the case studies focus on three different downstream classifications: political leaning detection, toxicity prediction, and sentiment classification. We end the paper in Section 6 with a discussion, report limitations and present avenues for future research.
@@ 2.1 Explainable AI and Counterfactual Explanations
As AI systems become increasingly integrated into business processes, the demand for transparency in these systems grows. In addition to the AI system being effective, it is often important that developers, users, and other stakeholders receive insight into the why and how behind the behavior of the system. These needs for transparency are described in detail by Martens and Provost (2014), while Meske et al. (2022) argue that Explainable AI (XAI) is a central issue for information systems research and should receive more attention.
The research field addressing techniques to provide such transparency has become known as Explainable AI (XAI). XAI techniques aim to make the internal mechanisms or the reasons for the input/output behavior of AI systems more comprehensible to humans (Gunning et al., 2019). More specifically, XAI methods examine a variety of different aspects of AI systems, and researchers and practitioners need to be precise about what sort of AI explanation they are talking about (Martens et al., 2025). The vast majority of XAI techniques focus on one of the following: building interpretable models, understanding the inner workings of complex models, identifying features that affect model scoring, or identifying features or parts of inputs that lead to a certain system behavior. This paper focuses on the last of those.
AI explanations serve a variety of purposes. They can help build user or management trust, assist model developers in debugging or model improvement, uncover hidden biases or fairness concerns, satisfy regulatory or legal requirements for transparency, improve relations with customers affected by AI systems actions, impact human-AI collaboration, etc. (Ferrario and Loi, 2022; Goethals et al., 2024; Martens and Provost, 2014; Vermeire et al., 2022; Wachter et al., 2017; von Zahn et al., 2025). XAI methods have been developed to explain both global behavior (how a model behaves across the input space), and local behavior (why a model made a specific prediction or decision for a given instance) (Martens and Provost, 2014; Guidotti et al., 2018; Molnar, 2020).
This paper focuses on explanations for local (instance-specific) input/output behavior of generative AI systems. We adapt a well-known instance-level explanation technique, namely counterfactual explanations (Fernández-Loría et al., 2022; Martens and Provost, 2014; Wachter et al., 2017). Counterfactual explanations aim to identify (minimal) changes to an input instance that would result in a different decision outcome. For example, in the context of credit scoring, a possible counterfactual explanation could be ‘If the loan amount had been \$16,000 lower, you would have received the loan’ (Fernández-Loría et al., 2022).
@@ 2.2 Explanations for Generative AI systems
LLMs like ChatGPT have been criticized for being opaque, black boxes (Das et al., 2025; Zhao et al., 2024). The internal workings of LLMs remain largely inscrutable, even to their developers, and foundational model providers rarely provide explanations for their systems’ behavior. This is particularly problematic given the systems’ remarkable performance and attraction to client organizations, as their opaqueness poses challenges for safety, accountability, human trust, and even debugging functional systems built based on foundational models. Explainable AI techniques are essential tools to foster a deeper understanding of LLM behavior and to mitigate the risks associated with their deployment (Das et al., 2025; Ferdaus et al., 2024).
@@
A significant portion of the literature on LLM interpretability focuses on model-centric approaches, specifically, approaches that reference the internals of the model. For example, representation surgery involves directly intervening in a model’s latent space to alter internal representations or isolate specific features (Ravfogel et al., 2025). More specifically, Ravfogel et al. (2025) study the idea of counterfactuals in the representation space, raising the question: What would the model have generated if its internal representation had undergone a specific intervention? [^4] As noted above, the term “counterfactual” is overloaded even in the XAI space; Ravfogel et al. (2025) employ a different notion of a counterfactual than the notion we adapt from the XAI literature. (To our knowledge the latter usage was the original use of the term for XAI.) Other methods, such as circuit analysis, attempt to decompose computations into interpretable submodules or “circuits” to trace how information flows through the network (Tang et al., 2023). Attention-based explanations focus on analyzing the knowledge encoded in attention weights (Zhao et al., 2024). These internalist approaches require deep technical expertise, are difficult to transfer across different model architectures, and may or may not produce an “explanation” for the AI system’s behavior that is actually interpretable to stakeholders.
@@
Our work fits into the body of research that adopts a prompt-centric perspective to analyzing LLM behavior, examining how variation in inputs, rather than model internals or training data, shape the behavior of LLMs. Numerous studies show that LLMs are highly prompt-sensitive, with even minor changes significantly affecting outputs (Anagnostidis and Bulian, 2024; Elazar et al., 2021; Rauba et al., 2024). Rauba et al. (2024) highlight the obstacle of disentangling meaningful changes in the output from the inherent stochasticity of LLM outputs, and propose a statistical framework to reformulate LLM perturbation analysis as a frequentist hypothesis problem. Mohammadi (2024) leverage Shapley values to expose the “token noise” effect, where seemingly unimportant tokens exert a disproportionate influence on model output. Hackmann et al. (2024) study the statistical impact of different words on the model output. Tools like Polyjuice generate diverse, plausible counterfactual prompts to support this line of inquiry (Wu et al., 2021). Other studies investigate the performance of using LLMs to generate counterfactual explanations themselves (Li et al., 2024; Youssef et al., 2024; Mayne et al., 2025).
@@
Our work examines how counterfactual explanations can be meaningfully applied to non-deterministic generative AI systems, when the focal characteristics of generative outputs can be revealed using downstream classification. To the best of our knowledge, we are the first to apply counterfactual explanations to full text outputs (continuations), whereas earlier work primarily focused on providing counterfactual explanations in contexts where generative AI systems were used to make explicit decisions, for example labeling (Chittimalla and Potluri, 2025; Mayne et al., 2025; Sarkar et al., 2024). [^5] An early short workshop paper on the present work appeared in 2025. That paper presented the general idea and some preliminary experiments. The current paper includes much more comprehensive scholarly discussion, an improved algorithm, and all new empirical results. We do not reference the paper here to maintain anonymity; however, we do cite it in the submission cover letter.
@@ 3 Problem formulation
Explainable AI can be confusing because there are so many different ways explanations might be defined. Here we follow and adapt previous work on counterfactual explanations for AI decisions (Fernández-Loría et al., 2022). However, there are several non-trivial challenges to doing so.
@@ 3.1 Four key challenges
    The majority of existing counterfactual methods assume feature-set input, and the methods operate on these sets of features; in particular, the ordering of the input is ignored. Although the original counterfactual explanation method operated on text as input (Martens and Provost, 2014), it essentially assumed a token-set representation (bag-of-words, tf-idf, etc.), without considering the sequential and linguistic structure of the text. While a token-set explanation method seems like a great place to start, we would not want to use it to define explanations for LLM-based systems, as the linguistic structure of the text may be critical. An exception to simple feature-set input in prior work is CFEs for image classification, where “pieces” (e.g., regions) of the image can be defined in different ways and then operated on to produce counterfactuals (Vermeire et al., 2022); we use this idea as inspiration for our adaptation to LLMs.
@@
    The set of counterfactual explanations that will be produced depends on how one searches through the space of collections of “pieces” of the input to remove (see above), and with what (if anything) one replaces them. How this replacement works is a crucial aspect of formulating explanations that will be effective for a particular problem (Fernández-Loría et al., 2022), and as Fernández et al. explained, the choice of replacement should depend on the specifics of the particular explanation context. Explanation methods should be flexible enough to allow for different choices. For LLM-based generative systems, we can envision many different replacement options, including dropping pieces of the prompt, masking them, replacing them with synonymous phrases, and so on. In addition, the structure of the text may lead to more intelligent search strategies for choosing combinations of “pieces” of the input to replace, which could also lead to significant speed-ups for the algorithm. A straightforward example would be to prefer combining adjacent input pieces (sequential tokens, phrases, sentences). One can envision more sophisticated techniques taking into account the vast amount of syntactic and semantic analysis provided by (computational) linguistics.
@@
    Finally, unlike prior applications of CF explanations, generative AI systems do not produce a fixed, deterministic output from a given input (Atil et al., 2024). This substantially complicates matters. CFEs ask some variant of “what can I change in the input to get a focal change in the output?” But just running the system again with the same input might produce the focal change in the output! Therefore, the question needs to be generalized to be probabilistic or otherwise distributional. For example, a PCE algorithm could ask about (changes to) the probability that a particular input will produce an output with/without the focal characteristic.
@@
In order to change the input, we mainly experiment with a straightforward approach that involves replacing the chosen segments of the prompt with a textual “masking” token (underscores). [^7] Using explicit token masking produced comparable results. Masking is different from simple token removal, because the masking tokens indicate to the LLM that there is in fact input there (it just is hidden). There also is a difference between the “replacement” strategy used to generate explanations, and a replacement strategy which may be applied subsequently in certain circumstances to produce suitable output, based on what the explanations revealed. The PCEs are an analysis tool. There always will still be the question: what should we do based on the results? [^8] The counterfactual explanations literature sometimes confounds the questions of what in its input caused the system to give a particular output—the primary CFE question—and what can be done in the real world based on the CFE answer. In the case studies, when looking into what we might do based on a PCE, we demonstrate alternatives to simply removing segments, such as replacing them with paraphrases when that makes sense for the application.
@@
To illustrate, let us say that the characteristic we are interested in is negative sentiment in our LLM’s generated output (in this case generated with GPT-2). Imagine that the generative system receives the prompt “Create a horrible short story about a dog” and our characteristic classifier is a sentiment classifier that classifies the polarity and magnitude of the sentiment of the generated output. Due to the word “horrible,” it would be natural to expect that the output for this prompt would have negative sentiment polarity. Figure 2(a) shows the distribution over the sentiment of the outputs over 100 different generations based on this one prompt. The outputs are often, but not always, classified as having negative sentiment.
(a) Sentiment score distribution over the outputs of the generative AI (GPT-2) after the prompt “Create a horrible short story about a dog” was run 100 times. We see that the sentiment of the generated output varies widely, although it very often is very negative. The vertical dashed line shows the average of the sentiment scores across the 100 runs.
(b) Sentiment score distribution over the outputs of the generative AI system (GPT-2) after the prompt “Create a short story about a dog” was run 100 times. We see that removing the word “horrible,” identified by the PCE, shifts the sentiment to be generally more positive (the dashed line shows the average sentiment score), and in particular shifts it to be above zero on average. Notably, now there are fewer cases where the generated outputs have very negative sentiment.
@@
Of course there is not a single right answer to either of these questions; we need to define what we care about for the particular task at hand. [^10] This is not a new observation; it is consistent with prior literature on counterfactual explanations for AI systems Fernández-Loría et al. (2022). Thus, a framework for creating PCEs for generative AI systems needs to be flexible. Do we care about the mean prediction probability of negative sentiment over all the samples? The median? Where do we want to set the threshold in any of these cases in order to say that the behavior has changed? How many runs should we look at? The definition of a PCE will depend on the user’s needs. For example, in one setting, we might care that the sentiment the system produces is generally positive. In another setting, we might be extremely averse to negative sentiment, and therefore specify that the likelihood of negative sentiment must be very small.
To illustrate, let’s say we decide to use the average (mean) of the characteristic scores (from the downstream classifier) across the empirical distribution, and we want this average to be positive. Based on this choice, we can define a PCE as a minimal set of “parts” of the input prompt that, if they were not present, would move the average of the prediction distribution from below to above the threshold (in this case 0). Our explanation algorithm (explained in detail in Section 4) for the dog story example leads to the not-surprising explanation: “If the word horrible were not part of the input prompt, the average sentiment of the generated output would be positive.” The resultant prediction distribution, with “horrible” removed, is also depicted in Figure 2(b). Importantly, we should note how the posing of the explanation question affects the result. Removing ”horrible” brings the average sentiment above zero, as requested. However, the LLM still generates a good number of negative-sentiment stories even with the seemingly innocuous prompt, ”Create a short story about a dog.” If the explanation requirement had been different, for example have a very low probability of very negative sentiment, we would have gotten a different (or no) explanation.
Importantly, as mentioned above, this tells us what is it about the prompt that resulted in the sentiment of the output to be negative. It does not (necessarily) tell us what to do about it. That is a different problem from creating the PCE. It is unclear for this example whether the system could create a horrible story about a dog that did not have consistently negative sentiment. In our case studies, we illustrate several possible ways to act on this information, such as by paraphrasing the “offending” parts of the input identified by the PCE.